# Testing Most Influential Sets

### v2

@@ Abstract
Small subsets of data with disproportionate influence on model outcomes can have dramatic impacts on conclusions, with a few data points sometimes overturning key findings. While recent work has developed methods to identify these most influential sets, no formal theory exists to determine when their influence reflects genuine problems rather than natural sampling variation. We address this gap by developing a principled framework for assessing the statistical significance of most influential sets. Our theoretical results characterize the extreme value distributions of maximal influence and enable rigorous hypothesis tests for excessive influence, replacing current ad-hoc sensitivity checks. We demonstrate the practical value of our approach through applications across economics, biology, and machine learning benchmarks.
@@ 1 Introduction
Machine learning (ML) models and statistical inferences can be highly sensitive to small subsets of data. In many applications, just a handful of samples can overturn key conclusions: two countries nullify the estimated effect of geography on development (Kuschnig et al., 2021), a single outlier flips the sign of a treatment effect (Broderick et al., 2021), or a small group of individuals drives disparate outcomes in algorithmic decision-making (Black & Fredrikson, 2021). These most influential sets — data subsets with the greatest influence on model predictions — are central to questions of interpretability, fairness, and robustness in modern machine learning (see, e.g., Black & Fredrikson, 2021; Chen et al., 2018; Chhabra et al., 2023; Ghorbani & Zou, 2019; Sattigeri et al., 2022).
Despite their practical importance, practitioners lack principled tools to assess whether a set’s influence is genuinely problematic. Current practice relies on domain expertise and ad-hoc sensitivity checks, while approximate methods such as influence functions (Koh & Liang, 2017; Fisher et al., 2023; Schioppa et al., 2023) systematically underestimate the impacts of sets and extreme cases (Basu et al., 2020; Koh et al., 2019). Recent work highlights both the promise and challenges of most influential subsets — small sets can drive results even in randomized trials (Broderick et al., 2021; Kuschnig et al., 2021), heuristic algorithms can fail in simple settings (Hu et al., 2024; Huang et al., 2025), and influence bounds remain an active area of research (Moitra & Rohatgi, 2023; Freund & Hopkins, 2023; Rubinstein & Hopkins, 2024). What remains missing is a principled method to distinguish natural sampling variation from genuinely excessive influence.
We develop a statistical framework for assessing the significance of most influential sets. By focusing on linear regression — a tractable, interpretable, and widely-used setting that underlies many modern methods (Rudin, 2019) — we derive the exact asymptotic distributions of maximal influence. We show that two distinct regimes emerge depending on the size of the influential set: when the size is fixed, maximal influence converges to a heavy-tailed Fréchet distribution; when the size grows with the sample, maximal influence converges to a well-behaved Gumbel distribution. Our results enable principled hypothesis tests for excessive influence, replacing ad-hoc diagnostics with rigorous statistical procedures. We demonstrate their practical value via applications across economics, biology, and machine learning benchmarks, resolving ambiguous cases where influential sets drive contested findings.
1.1 Contributions
We present a comprehensive analysis of the influence of most influential sets, both theoretically and in practical applications. To summarize, our main contributions are:
@@ 1.1 Contributions
    Efficient implementation. We provide computationally efficient procedures for evaluating influence, making our approach practical for real-world applications.
@@
    Empirical validation. We demonstrate the utility of our framework across domains, resolving the contested “Blessing of Bad Geography” in economics, assessing robustness in biological data of sparrow morphology, and auditing fairness in ML benchmark datasets.
1.2 Outline
@@ 2 Preliminaries and Background
  • Fairness auditing: An algorithmic decision-making system produces vastly different outcomes for a protected group. What if this disparity is explained by only a handful of data points?
  • Data cleaning: A single influential data point among a thousand samples flips a strong correlation to a null result. Should we trust the original finding or the one without the outlier?
@@ 2.1 Formal Problem Statement
Let $\left[N\right]=\left\{1,\dots,N\right\}$ and denote both an index set and its corresponding subsample as ${\mathbb{S}}\subset\left[N\right].$ For any subset $\mathbb{S}$}, we use a subscript $\hat{\theta}_{-{\mathbb{S}}}$ to denote a quantity $\theta$ without $\mathbb{S}$}, i.e.
@@
For a positive integer $k\ll N,$ the $k$-most influential subset is
${\mathbb{S}}_{k}^{\max}\coloneqq\underset{{\mathbb{S}}\subset\left[N\right],\lvert{\mathbb{S}}\rvert\leqslant k}{\arg\max}\Delta\left({\mathbb{S}};\phi\right),$
where $\Delta\left({\mathbb{S}};\phi\right)=\phi(\hat{\theta})-\phi(\hat{\theta}_{-{\mathbb{S}}})$ is the influence of subset $\mathbb{S}$} on target function $\phi:\mathbb{R}^{Q}\mapsto\mathbb{R}.$ The maximum influence is denoted as $\Delta^{\max}=\Delta\left({\mathbb{S}}_{k}^{\max};\phi\right).$
@@
What is the probability distribution of $\Delta^{\max},$ and how can we use it to distinguish excessive influence from natural sampling variation?
@@
A common approach to study influence is via influence functions (Fisher et al., 2023). These are motivated by reweighing via the perturbation
$\hat{\theta}(\epsilon;{\mathbb{S}})\coloneqq\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\frac{1}{N}\sum_{n\not\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right)+\epsilon\sum_{i\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{i}),y_{i}\right).$
Setting $\epsilon=0$ recovers $\hat{\theta}$, while $\epsilon=-N^{-1}$ yields $\hat{\theta}_{-{\mathbb{S}}}.$ The influence function is the linear approximation at $\epsilon=0$:
$\hat{\theta}(\epsilon;{\mathbb{S}})\approx\mathcal{I}({\mathbb{S}})\coloneqq\frac{\mathrm{d}\hat{\theta}(\epsilon;{\mathbb{S}})}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}.$ (1)
This yields the following first-order estimate of influence:
$\Delta\left({\mathbb{S}};\phi\right)\approx-N^{-1}\frac{\mathrm{d}\phi\left(\hat{\theta}(\epsilon;{\mathbb{S}})\right)}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}.$
@@
Exact Influence
In this work, we focus on the exact influence of subsets in a tractable but ubiquitous setting. We avoid linear approximations to accurately assess the most influential sets of interest, where extreme behavior dominates and first-order approximations fail most dramatically.
@@ 3 Proposed Approach
Consider the standard linear regression model, where $f$ is a linear function relating features $X\in\mathbb{R}^{P}$ to the outcome $Y\in\mathbb{R}.$ Stacking the observed training sample yields the design matrix $\mathbf{X}\in\mathbb{R}^{N\times P}$ and outcome vector $\mathbf{y}\in\mathbb{R}^{N}.$ We assume that $\mathbf{X}^{\prime}\mathbf{X}$ is invertible and remains so after removing any subset. [^1] Alternatively, one may consider the penalized estimator with $\mathbf{X}^{\prime}\mathbf{X}+\lambda\mathbf{I},$ where $\lambda>0$ creates a ridge that guarantees invertibility. Let $\mathbf{H}=\mathbf{X}\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}$ denote the hat matrix, with leverage scores along its diagonal $\mathbf{h}$. The ordinary least squares (OLS) estimator is
$\hat{\theta}=\underset{\theta}{\arg\min}\lVert Y-X\theta\rVert^{2}=\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}\mathbf{y},$
with predictions $\hat{\mathbf{y}}=\mathbf{X}\hat{\theta}$ and residuals $\mathbf{r}=\mathbf{y}-\hat{\mathbf{y}}$.
We focus on the influence of subsets on a particular coefficient of interest for interpretation. For simplicity, we assume a univariate model with a positive coefficient, orthogonalize features where necessary, and set the target function to $\phi(\theta)=\theta_{1}.$
3.1 Exact Influence Formulas
The influence of a single observation $i$ is well-known (Belsley et al., 1980; Cook, 1979) to be
$\Delta\left(\{i\}\right)=\frac{x_{i}r_{i}}{\sum_{n}x_{n}^{2}}\frac{1}{1-h_{i}},$ (2)
@@ 3.1 Exact Influence Formulas
For sets of observations, we can derive a closed-form expression via recursive application of Equation 2 and updating formulae for $r_{i}$ and $h_{i}.$ (Details are provided in the Appendix.) The following closed form solution is particularly convenient:
@@
The influence of set $\mathbb{S}$} on $\hat{\theta}$ is
$\Delta\left({\mathbb{S}}\right)=\frac{\sum_{i\in{\mathbb{S}}}x_{i}r_{i}}{\sum_{n\not\in{\mathbb{S}}}x_{n}^{2}}.$ (3)
Proof sketch.
For a single observation, one can show that $\Delta\left(i\right)=\frac{x_{i}r_{i}}{\sum_{n\neq i}x_{n}^{2}}.$ Then, let ${\mathbb{S}}=\left\{1,2\right\}$ and define $D\coloneqq\sum_{n}x_{n}^{2}$ for simplicity. Evidently,
$\Delta\left(\{1,2\}\right)$ $=\frac{x_{1}r_{1}}{D_{-\{1,2\}}+x_{2}^{2}}+\frac{x_{2}\left(r_{2}+x_{2}\Delta\left(\{1\}\right)\right)}{D_{-\{1,2\}}}$
$=\frac{x_{1}r_{1}}{D_{-\{1,2\}}+x_{2}^{2}}+\frac{x_{2}r_{2}}{D_{-\{1,2\}}}+\frac{x_{2}^{2}\left(x_{1}r_{1}\right)}{D_{-\{1,2\}}\left(D_{-\{1,2\}}+x_{2}^{2}\right)}$
$=\frac{\left(D_{-\{1,2\}}+x_{2}^{2}\right)x_{1}r_{1}}{D_{-\{1,2\}}\left(D_{-\{1,2\}}+x_{2}^{2}\right)}+\frac{x_{2}r_{2}}{D_{-\{1,2\}}}=\frac{x_{1}r_{1}+x_{2}r_{2}}{D_{-\{1,2\}}},$
where the second term in the first line corrects $r_{2}$ to reflect the removal of observation $1,$ which the second line expands. The third line merges terms one and three by transforming to a common denominator, and the fourth line simplifies the expression. Assuming this identity holds for $\lvert{\mathbb{S}}\rvert=K$ we can show by induction that it holds for $\lvert{\mathbb{S}}\rvert=K+1,$ and the result follows. Details are provided in the Appendix. ∎
Proposition 1 elegantly reveals the additive structure of individual contributions in the numerator and the multiplicative adjustment from the remaining data in the denominator. This representation enables efficient computation without explicitly forming leverage scores for each subset, making our approach computationally tractable for large datasets.
3.2 Extreme value distribution
We now turn to the distribution of $\Delta({\mathbb{S}})$ for the most influential set, ${\mathbb{S}}_{k}^{\max}.$ Since this quantity is defined by an extremal operation (maximization over all possible subsets), its asymptotic behavior is governed by extreme value theory. Specifically, we seek the limiting extreme value distribution (EVD) $H$ such that $\Delta^{\max}\in\mathrm{MDA}\left(H\right),$ i.e., $\Delta^{\max}$ lies in the maximum domain of attraction of $H.$

### v3

@@ Abstract
Small influential data subsets can dramatically impact model conclusions, with a few data points overturning key findings. While recent work identifies these most influential sets, there is no formal way to tell when maximum influence is excessive rather than expected under natural random sampling variation. We address this gap by developing a principled framework for most influential sets. Focusing on linear least-squares, we derive a convenient exact influence formula and identify the extreme value distributions of maximal influence — the heavy-tailed Fréchet for constant-size sets and heavy-tailed data, and the well-behaved Gumbel for growing sets or light tails. This allows us to conduct rigorous hypothesis tests for excessive influence. We demonstrate through applications across economics, biology, and machine learning benchmarks, resolving contested findings and replacing ad-hoc heuristics with rigorous inference.
@@ 1 Introduction
Machine learning (ML) models and statistical inferences can be highly sensitive to small subsets of data. In many applications, just a handful of samples can overturn key conclusions: two countries nullify the estimated effect of geography on development (Kuschnig et al., 2021), a single outlier flips the sign of a treatment effect (Broderick et al., 2023), or a small group of individuals drives disparate outcomes in algorithmic decision-making (Black and Fredrikson, 2021). These most influential sets — data subsets with the greatest influence on model predictions — are central to questions of interprtability, fairness, and robustness in modern machine learning (see, e.g., Black and Fredrikson, 2021; Chen et al., 2018; Chhabra et al., 2023; Ghorbani and Zou, 2019; Sattigeri et al., 2022).
Despite their practical importance, practitioners lack principled tools to assess whether a set’s influence is genuinely problematic. Current practice relies on heuristics, ad-hoc sensitivity checks, and domain expertise, while approximate methods such as influence functions (Koh and Liang, 2017; Fisher et al., 2023; Schioppa et al., 2023) systematically underestimate the impacts of sets and extreme cases (Basu et al., 2020; Koh et al., 2019). Recent work highlights both the promises and challenges of most influential subsets — small sets can drive results even in randomized trials (Broderick et al., 2023; Kuschnig et al., 2021), heuristic algorithms can fail in simple settings (Hu et al., 2024; Huang et al., 2025), and influence bounds remain an active area of research (Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Rubinstein and Hopkins, 2025). What remains missing is a principled method to distinguish natural sampling variation from genuinely excessive influence.
We develop a statistical framework for assessing the significance of most influential sets. By focusing on linear regression — a tractable, interpretable, and widely-used setting that underlies many modern methods — we derive the exact asymptotic distributions of maximal influence. We show that two distinct regimes emerge depending on the size of the influential set: when the size is fixed, maximal influence converges to a heavy-tailed Fréchet distribution; when the size grows with the sample, maximal influence converges to a well-behaved Gumbel distribution. Our results enable principled hypothesis tests for excessive influence, replacing ad-hoc diagnostics with rigorous statistical procedures. We demonstrate their practical value via applications across economics, biology, and machine learning benchmarks, resolving ambiguous cases where influential sets drive contested findings.
Contributions.
We present a comprehensive analysis of the influence of most influential sets, both theoretically and in practical applications. Our main contributions are:
@@
    Efficient implementation. We provide a closed-form formula for evaluating set influence, making our approach practical for real-world applications.
@@
    Empirical validation. We demonstrate the utility of our framework across domains, resolving the contested ‘Blessing of Bad Geography’ in economics, assessing robustness in biological data of sparrow morphology, and auditing fairness in ML benchmark datasets.
To summarize — we provide the first rigorous theoretical results that allow us to interpret influence, and demonstrate practicality by resolving contested findings.
Outline.
@@ 2 Preliminaries and Background
  • Fairness auditing: An algorithmic decision-making system produces different outcomes for a protected group. What if the disparity can be explained by only a handful of data points?
  • Data cleaning: A single influential point among a thousand samples flips a strong correlation to a null result. Should we trust the original finding or the one without the outlier?
@@
What has been unclear, however, is how to interpret and deal with influential sets. Existing influence methods quantify how much changes, but offer little guidance on interpretation. Current practice relies heavily on domain expertise and ad-hoc rules, lacking a statistically rigorous framework for judging influence. Heuristics, such as sign-flips or significance thresholds, can flag effects that are stable and miss genuinely problematic influence. We address this gap by quantifying whether observed maximum influence is statistically compatible with natural sampling variation.
@@ 2.1 Formal Problem Statement
Let $\left[N\right]=\left\{1,\dots,N\right\}$ and denote both an index set and its corresponding subsample as ${\mathbb{S}}\subset\left[N\right].$ For any subset $\mathbb{S}$}, we use a subscript $\hat{\theta}_{-{\mathbb{S}}}$ to denote an estimate $\hat{\theta}$ without $\mathbb{S}$}, i.e.
@@
For a positive integer $k\ll N,$ the $k$-most influential set is
${\mathbb{S}}_{k}^{\max}\coloneqq\underset{{\mathbb{S}}\subset\left[N\right],\lvert{\mathbb{S}}\rvert=k}{\arg\max}\Delta\left({\mathbb{S}};\phi\right),$
where $\Delta\left({\mathbb{S}};\phi\right)=\phi(\hat{\theta})-\phi(\hat{\theta}_{-{\mathbb{S}}})$ is the influence of subset $\mathbb{S}$} on the scalar target function $\phi:\mathbb{R}^{Q}\mapsto\mathbb{R}.$ We denote the maximum influence as $\Delta^{\max}=\Delta\left({\mathbb{S}}_{k}^{\max};\phi\right).$
@@
What is the probability distribution of $\Delta^{\max},$ and how can we distinguish excessive influence from natural sampling variation?
@@
A common and related approach to study influence is via influence functions (Fisher et al., 2023; Hu et al., 2024; Koh and Liang, 2017). These are motivated by reweighing via the perturbation
$\hat{\theta}(\epsilon;{\mathbb{S}})\coloneqq\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\frac{1}{N}\sum_{n=1}^{N}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right)+\epsilon\sum_{i\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{i}),y_{i}\right).$
Setting $\epsilon=0$ recovers $\hat{\theta}$, while $\epsilon=-N^{-1}$ yields $\hat{\theta}_{-{\mathbb{S}}}.$ The influence function is the first-order linear approximation at $\epsilon=0.$
@@
Exact Maximum Influence.
We therefore derive an exact influence formula (in Section 3) and characterize the behavior of maximum influence (Section 3.1) to enable principled testing (Section 3.3). We focus on the tractable but ubiquitous linear setting, allowing us to accurately portray most influential sets, for which extreme behavior dominates and first-order approximations fail most dramatically.
2.2.1 Extreme Value Theory
Our goal is to characterize the behavior of $\Delta^{\max},$ the influence of the most influential set. This quantity is defined through maximization over all possible subsets, and its distribution is governed by extreme value theory rather than classical asymptotics. See Figure 1 for an illustration. When taking maxima over random quantities, three possible limiting distributions can emerge (Fisher and Tippett, 1928; Gnedenko, 1943): the well-behaved Gumbel (Type I), the heavy-tailed Fréchet (Type II), and the bounded Weibull (Type III).
Figure 1: Illustration of the distribution of $\Delta(\{i\})$ for general observations, and the maximum influence $\Delta^{\max}$ over all possible $i,$ for $N\in\{25,50,100\}.$ One can clearly discern an upward shift and a substantial increase of the density in the tails.
We need to determine which extreme value distribution (EVD) attracts $\Delta^{\max}.$ Specifically, we distinguish between the Gumbel distribution with exponential tails, and the Fréchet distribution with polynomial tails allowing for potentially arbitrarily large influence. (The Weibull distribution can be ruled out since influence is unbounded.) Once we establish the asymptotic distribution, we show how it applies to finite samples.
2.3 Setting
Consider the standard linear regression model, where $f$ is a linear function relating random features $X$ to the outcome $Y$ via the parameter vector $\theta$. The ordinary least-squares (OLS) population estimator of $\theta$ is
$\tilde{\theta}\coloneqq\mathbb{E}[XX^{\prime}]^{-1}\mathbb{E}[XY]$
yielding fitted values $\hat{Y}=X^{\prime}\tilde{\theta}$ and the associated residuals $R=Y-X^{\prime}\tilde{\theta}$. Stacking $N$ observed training samples yields the design matrix $\mathbf{X}\in\mathbb{R}^{N\times P}$ and outcome vector $\mathbf{y}\in\mathbb{R}^{N}.$ For OLS, we assume that $\mathbf{X}^{\prime}\mathbf{X}$ is invertible and remains so after removing any subset. [^1] This is not necessary for our results, as they extend to ridge regression, where the penalization parameter $\lambda>0$ in $\mathbf{X}^{\prime}\mathbf{X}+\lambda\mathbf{I}$ creates a ridge that guarantees invertibility. The sample OLS estimator is then
$\hat{\theta}\coloneqq\underset{\theta}{\arg\min}\lVert\mathbf{y}-\mathbf{X}\theta\rVert^{2}=\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}\mathbf{y},$
yielding predictions $\hat{\mathbf{y}}=\mathbf{X}\hat{\theta}$ and residuals $\mathbf{r}=\mathbf{y}-\hat{\mathbf{y}}$.
For our main results, we consider least-squares estimates of the regression coefficients, with and without penalization. For illustration, however, we will consider a univariate model with one positive coefficient of interest and target function $\phi(\theta)=\theta_{1}.$ This eases notation considerably and comes without loss of generality — theoretical results apply to more general settings. [^2] Multiple features can be factored out via the Frisch-Waugh-Lovell theorem under mild assumptions, leaving an equivalent univariate regression. A different sign can be accommodated by simply sign-flipping the feature of interest or adjusting the target function $\phi$ accordingly.
@@ 3 Proposed Approach
The influence of a single observation $i$ is well-known (Belsley et al., 1980; Cook, 1979; Walker and Birch, 1988) to be
$\Delta\left(\{i\}\right)=\frac{\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{x}_{i}r_{i}}{1-h_{i}},$
@@
We extend this result to be (i) applicable to sets of observations, and (ii) analytically convenient.
@@
The influence of some set $\mathbb{S}$} on the least-squares estimator $\hat{\theta}$ is
$\Delta\left({\mathbb{S}}\right)=\left(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I}_{P}\right)^{-1}\mathbf{X}_{{\mathbb{S}}}^{\prime}\mathbf{r}_{{\mathbb{S}}},$ (1)
where $\lambda\geqslant 0$ is an optional penalization parameter.
Proof.
Partition the full normal equations into $\mathbb{S}$} and $\mathbb{S}$}, and subtract $\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{X}_{\mathbb{S}}\hat{\theta}$:
$(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})\,\hat{\theta}$ $=\mathbf{X}^{\prime}\mathbf{y}$
$(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{X}_{\mathbb{S}}+\lambda\mathbf{I})\,\hat{\theta}$ $=\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{y}_{-{\mathbb{S}}}+\mathbf{X}_{{\mathbb{S}}}^{\prime}\mathbf{y}_{\mathbb{S}}$
[…]
The key element stems from $\mathbf{r}_{\mathbb{S}}=\mathbf{y}_{\mathbb{S}}-\mathbf{X}_{\mathbb{S}}\,\hat{\theta}$, and the result follows from subtracting the leave-out normal equations and inverting:
$(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I})\,\hat{\theta}$ $=\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{y}_{-{\mathbb{S}}}+\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{r}_{\mathbb{S}}$
$(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I})\,\left(\hat{\theta}-\hat{\theta}_{-{\mathbb{S}}}\right)$ $=\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{r}_{\mathbb{S}}$
$\Delta({\mathbb{S}})$ $=(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I})^{-1}\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{r}_{\mathbb{S}}.$
We thank an anonymous reviewer for pointing us towards this elegant proof. ∎
Proposition 1 reveals the additive structure of individual contributions in the numerator, as well as the multiplicative adjustment from the denominator. It provides an exact closed-form expression that avoids re-fitting the model for each candidate subset.
3.1 Extreme Value Distributions
We now turn to finding the distribution of $\Delta({\mathbb{S}})$ for the most influential set, ${\mathbb{S}}_{k}^{\max}.$ Since this quantity is defined by an extremal operation (maximization over all possible subsets), its asymptotic behavior is governed by extreme value theory. Specifically, we seek the limiting EVD $H$ such that $\Delta^{\max}\in\mathrm{MDA}\left(H\right),$ i.e., $\Delta^{\max}$ lies in the maximum domain of attraction of $H.$